OpenStudy (anonymous):

a recipe calls for 4 4/5 tbsp of butter ` sarah is making 3 1/3 times the recipe for her soccer team ; how much butter does she need ? write your answer in simplest form .

OpenStudy (anonymous):

\[4\tfrac{4}{5}\times 1\tfrac{3}{3}\]

OpenStudy (anonymous):

change each mixed number to an improper fraction and then multiply

OpenStudy (anonymous):

\[4\tfrac{4}{5}=\frac{24}{5}\] \[3\tfrac{1}{3}=\frac{10}{3}\]

OpenStudy (anonymous):

\[\frac{24}{5}\times \frac{10}{3}=8\times 2 = 16\]

OpenStudy (anonymous):

because 5 goes into 10 two times and 3 goes into 24 eight times